# Data-adaptive structural change-point detection via isolation

### v2

@@ Appendix D Proof of Theorem 2
See Lemma 7 from [6].
The steps of the proof of Theorem 1 are the same as for Theorem 2.
@@
$\mathbb{P}\Biggl{(}\hat{N}=N,\max_{j=1,2,\ldots,N}\biggl{(}\left|\hat{r}_{j}-r_{j}\right|\left(\Delta^{f}_{j}\right)^{2/3}\biggl{)}\leq C_{3}(\log T)^{1/3}\Biggl{)}\geq 1-\frac{1}{6\sqrt{\pi}{T}},$ (48)
which implies result (13).
@@
$B_{T}^{\ast}=\left\{\max_{1\leq s\leq b<e<{T}}\max_{\begin{subarray}{c}r_{j-1}<s\leq r_{j}\\ r_{j}<e\leq r_{j+1}\\ s\leq b<e\end{subarray}}\frac{\left|\left\langle\boldsymbol{\phi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{b}}\rangle-\boldsymbol{\phi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{r}}\rangle,\boldsymbol{\epsilon}\right\rangle\right|}{\|\boldsymbol{\phi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{b}}\rangle-\boldsymbol{\phi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{r}}\rangle\|_{2}}\leq\sqrt{8\log T}\right\}.$ (49)
@@
where $C^{\ast}$ satisfies assumption (A3), $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{\log T}$ and $3/2\leq n\leq\delta_{T}/2\lambda_{T}$. As before, for $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). For $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\}$ being the location of the largest difference detected in the interval $[s,e]$, $1\leq s<e\leq T$, define $c_{m}^{l}$ and $c_{k}^{r}$ as in (C). Since the length of the intervals in (5) is $\delta_{T}/2n$ and $\lambda_{T}\leq\delta_{T}/2n$, for $m,k\in\{0,1,\ldots,K^{\max}\}$ we ensure that there exists at least one $m$ and at least one $k$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$.
@@
Without loss of generality, we suppose that the first change-point to get detected is $r_{J}$ for some $J\in\left\{1,2,\ldots,N\right\}$. We can show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that $r_{J}$ is isolated using exactly the same argument as in (C). We will now show that for $\tilde{b}_{J}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C^{t}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})$, then $C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})>\zeta_{T}$. Using (D), we have that
$C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})\geq C^{r_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})\geq C^{r_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{f})-\sqrt{8\log T}.$ (50)
@@
$\min\{c_{k}^{r}-r_{J},r_{J}-c_{m}^{l}\}\geq\frac{\delta_{T}}{2n}.$ (51)
follows the exact same steps as Step 3 in the proof of Theorem 1 and will not be repeated. Now, using assumption (A3) and the results in (50), (51), we have that
@@
$=\biggl{(}\frac{1}{8\sqrt{3n^{3}}}-\frac{2\sqrt{2\log T}}{\delta_{T}^{3/2}\underline{f}_{T}}\bigg{)}\delta_{T}^{3/2}\underline{f}_{T}\geq\biggl{(}\frac{1}{8\sqrt{3n^{3}}}-\frac{2\sqrt{2}}{C^{\ast}}\bigg{)}\delta_{T}^{3/2}\underline{f}_{T}$ (52)
@@
$\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})\Bigl{)}^{2}>\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})\Bigl{)}^{2}.$ (53)
Proving (53) and using the definition of $b_{J}$, we can conclude that $\lvert b_{J}-r_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2}\leq\tilde{\gamma}_{T}$. Now, using (9), it can be shown, for $\boldsymbol{\phi_{s,e}^{b}}$ as defined in (12), that (53) is equivalent to
@@
$+2\Big{\langle}\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{\ast}}}\langle\boldsymbol{f},\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{\ast}}}\rangle-\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J}}}\langle\boldsymbol{f},\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J}}}\rangle,\boldsymbol{\epsilon}\Big{\rangle}.$ (54)
@@
$\Lambda\colon=\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{f})\Bigl{)}^{2}-\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{f})\Bigl{)}^{2}.$ (55)
For the right-hand side of (D) using (D),
$\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}-\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}\leq\max_{s,e,b:s\leq b<e}\Bigl{(}C^{b}_{s,e}(\boldsymbol{\epsilon})\Bigl{)}^{2}-\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}\leq 8\log T$ (56)
Using Lemma 3, (D) and (55), we have that for the left-hand side,
@@
$=2\sqrt{\Lambda}\sqrt{8\log T}.$ (57)
Using (55), (56) and (D), we can conclude that (D) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution
@@
$\Leftrightarrow\min\{\lvert r_{J}-b^{\ast}\rvert,r_{J}-c_{m^{\ast}}^{l}\}>\frac{\left(63\log T\right)^{1/3}\left(2\sqrt{2}+4\right)^{2/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}=\frac{C_{3}\left(\log T\right)^{1/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}$ (58)
@@
$\min\{r_{J}-c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}-r_{J}\}>2^{1/3}C_{3}\frac{\left(\log T\right)^{1/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}-1,$ (59)
@@
and we can deduce that (D) is restricted to
@@
which implies (53). So, we conclude that necessarily
$\lvert r_{J}-b_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}.$ (60)
But (59) must be true since if we assume that
@@
Thus, we have proved that for $\lambda_{T}\leq\delta_{T}/2n$, working under the assumption that both $A_{T}^{\ast}$ and $B_{T}^{\ast}$ hold, there will be an interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$ with $C^{b_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})>\zeta_{T}$, where $b_{J}=\textrm{argmax}_{c_{m^{\ast}}^{l}\leq t<c_{k^{\ast}}^{r}}C^{t}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})$ is the estimated location for the change-point $r_{J}$ that satisfies (60).
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (C). Using Lemma 3, we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}C^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\left(\boldsymbol{X}\right)$,
@@
$\geq\frac{1}{\sqrt{24}}\left(\min\{r_{J-1}-c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}-r_{J-1}\}\right)^{3/2}\Delta^{f}_{J-1}-\sqrt{8\log T}.$ (61)
Before we show that $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}\}\geq\delta_{T}/2n$, we need show that $c_{m^{\ast}}^{l}$ satisfies $c_{m^{\ast}}^{l}-r_{J-1}\geq\delta_{T}/2n$ The proof is exactly the same as in Step 5 of the proof of Theorem 1. It can be deduced that the right end-point of the interval will satisfy $c_{k_{J-1}}^{r}-r_{J-1}>\delta_{T}/2n$ for some $k_{J-1}$. For the left end-point of the interval, the same holds as in Step 3. In any case, $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}\}>\delta_{T}/2n$ holds and so, from (D), using exactly the same calculations as (D), we have that
@@
Appendix E Outline of the proof of Theorem 3
[…]
Step 3: Working under the assumption that both $A^{\ast}_{T}$ (and thus also $A_{T}$) and $B_{T}$ hold, we show that detection in an interval where the change-point is isolated will occur and so the proof can be restricted to an interval with a single change-point and we also prove that the estimated and true change-points are close to each other. The values used in this step for the positive constants $C_{1},C_{2},$ and $C_{3}$ of Theorem 3, are $C_{1}=\sqrt{C_{3}}+\sqrt{8},C_{2}=\frac{1}{\sqrt{4n}}-\frac{2\sqrt{2}}{\underline{C}_{M}},C_{3}=2(2\sqrt{2}+4)^{2}$. The order with which the change-points are detected depends on the location of the largest difference at each step of the algorithm and so, without loss of generality, we assume that the first change-point to get detected is $r_{J}$ for $J\in\{1,\ldots,N\}$. For $c_{m}^{l},c_{k}^{r}$ as defined in (C), $c_{m}^{l}\in I_{J}^{L}$ or $c_{k}^{r}\in I_{J}^{R}$ hold (or both), with $I_{J}^{L},I_{J}^{R}$ as defined in (5), and showing that the change-point will be isolated in $[c_{m}^{l},c_{k}^{r}]$ is the same as for Theorem1 Step 3. It can be shown that for $\tilde{b}={\rm argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C_{c_{m}^{l},c_{k}^{r}}^{t}$, $C_{c_{m}^{l},c_{k}^{r}}^{\tilde{b}}>\zeta_{T,d}$, so $r_{J}$ can be detected in the interval.
Now, for $k^{\ast}\in\{1,\ldots,K^{\max}\}$ and $m^{\ast}\in\{k^{\ast}-1,k^{\ast}\}$, we denote by $c_{m^{\ast}}^{l}\geq c_{m}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{k}^{r}$ the first left- and right-expanding points, respectively, where detection of $r_{J}$ occurs. Showing that the estimated and true change-points are close to each other requires working on the two mean-dominant norms of (21) separately. Denoting by $q_{J}:={\rm argmax}_{j=1,\ldots,d}\left|\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b_{J},j}\right|$ and $\alpha=\begin{cases}\Delta^{q_{J}}_{J},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ L_{2}(\boldsymbol{\Delta_{J}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases}$, we show that $\left|b_{J}-r_{J}\right|\alpha^{2}\leq C_{3}\log\left(Td^{\frac{1}{4}}\right).$ This can be done using a contradiction argument, based on the fact that for any $b^{*}\in\left\{c_{m^{\ast}}^{l},\ldots,c_{k^{\ast}}^{r}-1\right\}$ with $\left|b^{*}-r_{J}\right|\alpha^{2}>\log\left(Td^{\frac{1}{4}}\right)$, it holds that $\left(\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J},q_{J}}\right)^{2}>\left(\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{*},q_{J}}\right)^{2}$, which, using the definition of $b_{J}$, cannot be true. The detailed proof of this is skipped, but we highlight that it basically relies on a combination of the strategy followed in [3] that the consistency of the MID algorithm was proved, and the approach followed in our Theorem 1, Step 3.
Step 4: We show that when the algorithm re-starts, no change-points are left undetected and that the new intervals allow for the detection of the remaining change-points. The first part is immediate, as from Step 3 we know that the change-point was isolated in the interval in which detection occurred. Since the change-point was detected in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, MDAIS restarts on $[1,c_{m^{\ast}}^{l}]$ and $[c_{k^{\ast}}^{r},T]$. The proof that detection of $r_{j}$ occurs is the same as in Step 3 for all $j\in\{1,\ldots,J-2,J+2,\ldots,N\}$. We focus on $j=J-1$ and similar steps hold for $j=J+1$. For $k_{J-1}\in\{1,\ldots,K^{\max}\}$ and $m_{J-1}\in\{k_{J-1}-1,k_{J-1}\}$, we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$. The upper bound of the last inequality is the reason why $r_{J-1}$ should be considered separately, as the right end-point of the interval where detection will occur is upper bounded by the left end-point of the interval in which $r_{J}$ was detected. The detection of $r_{J-1}$ is at $b_{J-1}={\rm argmax}_{c_{m^{\ast}_{J-1}}^{l}\leq t<c_{k^{\ast}_{J-1}}^{r}}C_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{t}$, which, for $q_{J-1}:={\rm argmax}_{j=1,2,\ldots,d}\left|\tilde{X}_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{b_{J-1},j}\right|$, satisfies
$C_{3}\log\left(Td^{1/4}\right)\geq\begin{cases}\left|b_{J-1}-r_{J-1}\right|\left(\Delta_{J-1}^{q_{J-1}}\right)^{2},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ \left|b_{J-1}-r_{J-1}\right|L_{2}^{2}(\boldsymbol{\Delta_{J-1}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases}.$
Following similar steps as in Step 3, we have that $C_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{b_{J-1}}>\zeta_{T,d}$. Now, with
$\alpha_{J-2}:=\begin{cases}\Delta_{J-1}^{q_{J-1}},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ L_{2}(\boldsymbol{\Delta_{J-1}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases},$
it is straightforward to show, following the same process as Step 3 of the proof, that $\left|b_{J-2}-r_{J-2}\right|\alpha_{J-2}^{2}\leq C_{3}\log\left(Td^{1/4}\right)$.
and the algorithm terminates.

### v3

@@ Appendix C Outline of the proof of Theorem 3
See Lemma 7 from [1]. ∎
The steps of the proof of Theorem 2 are the same as for Theorem 1, which can be found in Appendix B. We work under the notation in (12), (13) and (14).
@@
$\mathbb{P}\Biggl{(}\hat{N}=N,\max_{j=1,2,\ldots,N}\biggl{(}\left|\hat{r}_{j}-r_{j}\right|\left(\Delta^{f}_{j}\right)^{2/3}\biggl{)}\leq C_{3}(\log T)^{1/3}\Biggl{)}\geq 1-\frac{1}{6\sqrt{\pi}{T}},$ (S1)
which implies result (18) of the main paper.
@@
$B_{T}^{\ast}=\left\{\max_{j=1,2\ldots,N}\max_{\begin{subarray}{c}r_{j-1}<s\leq r_{j}\\ r_{j}<e\leq r_{j+1}\\ s\leq b<e\end{subarray}}\frac{\left|\left\langle\boldsymbol{\phi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{b}}\rangle-\boldsymbol{\phi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{r}}\rangle,\boldsymbol{\epsilon}\right\rangle\right|}{\|\boldsymbol{\phi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{b}}\rangle-\boldsymbol{\phi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\phi_{s,e}^{r}}\rangle\|_{2}}\leq\sqrt{8\log T}\right\}.$ (S2)
@@
where $C^{\ast}$ satisfies Assumption (A3), $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{\log T}$ and $n\geq 3/2$. As before, for $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). For $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\}$ being the location of the largest difference detected in the interval $[s,e]$, $1\leq s<e\leq T$, define $c_{m}^{l}$ and $c_{k}^{r}$ as in (B) of the main paper. Since the length of the intervals in (5) is $\delta_{T}/2n$, $\lambda_{T}\leq\delta_{T}/2n$ ensures that there exists at least one $m\in\{0,1,\ldots,K^{\max}\}$ and at least one $k\in\{1,2\ldots,K^{\max}\}$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$.
@@
Without loss of generality, we suppose that the first change-point to get detected is $r_{J}$ for some $J\in\left\{1,2,\ldots,N\right\}$. We can show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that $r_{J}$ is isolated using exactly the same argument as in (B7) of the main paper. We will now show that for $\tilde{b}_{J}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C^{t}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})$, then $C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})>\zeta_{T}$. Using (S3), we have that
$C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})\geq C^{r_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})\geq C^{r_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{f})-\sqrt{8\log T}.$ (S3)
@@
$\min\{c_{k}^{r}-r_{J},r_{J}-c_{m}^{l}\}\geq\frac{\delta_{T}}{2n}.$ (S4)
follows the exact same steps as Step 3 in the proof of Theorem 1 and will not be repeated. Now, using Assumption (A3) and the results in (S3), (S4), we have that
@@
$=\biggl{(}\frac{1}{8\sqrt{3n^{3}}}-\frac{2\sqrt{2\log T}}{\delta_{T}^{3/2}\underline{f}_{T}}\bigg{)}\delta_{T}^{3/2}\underline{f}_{T}\geq\biggl{(}\frac{1}{8\sqrt{3n^{3}}}-\frac{2\sqrt{2}}{C^{\ast}}\bigg{)}\delta_{T}^{3/2}\underline{f}_{T}$ (S5)
@@
$\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})\Bigl{)}^{2}>\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})\Bigl{)}^{2}.$ (S6)
Proving (S6) and using the definition of $b_{J}$, we can conclude that $\lvert b_{J}-r_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2}\leq\tilde{\gamma}_{T}$. Now, using (11), it can be shown, for $\boldsymbol{\phi_{s,e}^{b}}$ as defined in (17), that (S6) is equivalent to
@@
$+2\Big{\langle}\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{\ast}}}\langle\boldsymbol{f},\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{\ast}}}\rangle-\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J}}}\langle\boldsymbol{f},\boldsymbol{\phi_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J}}}\rangle,\boldsymbol{\epsilon}\Big{\rangle}.$ (S7)
@@
$\Lambda\colon=\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{f})\Bigl{)}^{2}-\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{f})\Bigl{)}^{2}.$ (S8)
For the right-hand side of (S3) using (S3),
$\Bigl{(}C^{b^{\ast}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}-\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}\leq\max_{s,e,b:s\leq b<e}\Bigl{(}C^{b}_{s,e}(\boldsymbol{\epsilon})\Bigl{)}^{2}-\Bigl{(}C^{r_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{\epsilon})\Bigl{)}^{2}\leq 8\log T.$ (S9)
Using Lemma 3, (S3) and (S8), we have that for the left-hand side,
@@
$=2\sqrt{\Lambda}\sqrt{8\log T}.$ (S10)
Using (S8), (S9) and (S3), we can conclude that (S3) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution
@@
$\Leftrightarrow\min\{\lvert r_{J}-b^{\ast}\rvert,r_{J}-c_{m^{\ast}}^{l}\}>\frac{\left(63\log T\right)^{1/3}\left(2\sqrt{2}+4\right)^{2/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}=\frac{C_{3}\left(\log T\right)^{1/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}$ (S11)
@@
$\min\{r_{J}-c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}-r_{J}\}>2^{1/3}C_{3}\frac{\left(\log T\right)^{1/3}}{\Bigl{(}\Delta_{J}^{f}\Bigr{)}^{2/3}}-1,$ (S12)
@@
and we can deduce that (S3) is restricted to
@@
which implies (S6). So, we conclude that necessarily
$\lvert r_{J}-b_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}.$ (S13)
But (S12) must be true since if we assume that
@@
Thus, we have proved that for $\lambda_{T}\leq\delta_{T}/2n$, working under the assumption that both $A_{T}^{\ast}$ and $B_{T}^{\ast}$ hold, there will be an interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$ with $C^{b_{J}}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})>\zeta_{T}$, where $b_{J}=\textrm{argmax}_{c_{m^{\ast}}^{l}\leq t<c_{k^{\ast}}^{r}}C^{t}_{c_{{m^{\ast}}}^{l},c_{{k^{\ast}}}^{r}}(\boldsymbol{X})$ is the estimated location for the change-point $r_{J}$ that satisfies (S13).
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (B) of the main paper. Using Lemma 3, we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}C^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\left(\boldsymbol{X}\right)$,
@@
$\geq\frac{1}{\sqrt{24}}\left(\min\{r_{J-1}-c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}-r_{J-1}\}\right)^{3/2}\Delta^{f}_{J-1}-\sqrt{8\log T}.$ (S14)
Before we show that $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}\}\geq\delta_{T}/2n$, we need show that $c_{m^{\ast}}^{l}$ satisfies $c_{m^{\ast}}^{l}-r_{J-1}\geq\delta_{T}/2n$ The proof is exactly the same as in Step 5 of the proof of Theorem 1. It can be deduced that the right end-point of the interval will satisfy $c_{k_{J-1}}^{r}-r_{J-1}>\delta_{T}/2n$ for some $k_{J-1}$. For the left end-point of the interval, the same holds as in Step 3. In any case, $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}\}>\delta_{T}/2n$ holds and so, from (S3), using exactly the same calculations as (S3), we have that